OpenStudy (anonymous):

Help Please! The vertices of a rectangle are R(-5,-5), S(-1,-5), T(-1,1), and U(-5,1). After translation, R’ is the point (-11,-11). Find the translation rule and coordinates of U’. (x, y) --> (x-6, y+6); (-11, 7) (x, y) --> (x-6, y-6); (-11, -5) (x, y) --> (x+6, y-6); (1, -5) (x, y) --> (x-6, y+6); (1, 7)

OpenStudy (anonymous):

-11=-5+x x=-11+5=-6 -11=-5+y y=-11+5=-6 \[translation is (x,y)\rightarrow \left( x-6,y-6 \right)\] U' is (-5-6,1-6) i.e.,(-11,-5)

OpenStudy (anonymous):

Thank you.